# QA-MDT: Quality-aware Masked Diffusion Transformer for Enhanced Music Generation

### v3

@@ Abstract
Text-to-music (TTM) generation, which converts textual descriptions into audio, opens up innovative avenues for multimedia creation. Achieving high quality and diversity in this process demands extensive, high-quality data, which are often scarce in available datasets. Most open-source datasets frequently suffer from issues like low-quality waveforms and low text-audio consistency, hindering the advancement of music generation models. To address these challenges, we propose a novel quality-aware training paradigm for generating high-quality, high-musicality music from large-scale, quality-imbalanced datasets. Additionally, by leveraging unique properties in the latent space of musical signals, we adapt and implement a masked diffusion transformer (MDT) model for the TTM task, showcasing its capacity for quality control and enhanced musicality. Furthermore, we introduce a three-stage caption refinement approach to address low-quality captions’ issue. Experiments show state-of-the-art (SOTA) performance on benchmark datasets including MusicCaps and the Song-Describer Dataset with both objective and subjective metrics. Demo audio samples are available at https://qa-mdt.github.io/, code and pretrained checkpoints are open-sourced at https://github.com/ivcylc/OpenMusic.
@@ 1 Introduction
Figure 1: (a) The distribution curves of CLAP similarity and pseudo-MOS for large-scale open-source music databases AudioSet (Defferrard et al., 2016) and FMA (Defferrard et al., 2016), where darker areas represent higher text-audio consistency or audio quality. (b) FAD scores on the MusicCaps test set obtained from models trained for 50K steps on AudioSet and FMA, using filter ratios of 0%, 33%, 66%, and 100% of low-quality data. Here, low-quality data is determined by a Pseudo-MOS score (Ragano et al., 2023) of less than 4.0. It can be inferred that performance consistently worsens with larger filter ratios.
@@
In the TTM field, high-quality music signals is scarce. This prevalent issue of low-quality data, highlighted in Figure 1 (a), manifests in two primary challenges. Firstly, most available music signals often suffer from distortion due to noise, low recording quality, or outdated recordings, resulting in diminished generated quality, as measured by pseudo-MOS scores from quality assessment models (Ragano et al., 2023). Secondly, there is a weak correlation between music signals and captions, characterized by missing, weak, or incorrect captions, leading to low text-audio similarity, which can be indicated by CLAP scores (Wu* et al., 2023). These challenges, especially the inherent quality in the music signal itself, significantly hinder the training of high-performance music generation models, resulting in poor rhythm, noise, and inconsistencies with textual control conditions in the generated audio. Additionally, as shown in Figure 1 (b), directly filtering low-quality music, which robustly reduces the dataset size, leads to a consistent decline in model performance. Therefore, finding an effective training strategy for large-scale datasets with low-quality waveforms, mismatches, and missing labels has become an urgent challenge.
@@ 2 Related Work
Previous research has made efforts to improve the quality of generated audio, particularly in two key areas: waveform fidelity and the consistency between input text and generated content. Waveform quality can be compromised by issues like aliasing from low sampling rates and limited expressiveness due to monophonic representations, while models like MusicGen (Copet et al., 2024) and Stable Audio (Evans et al., 2024a, b), which directly model 32k and 44.1k stereo audio, have significantly enhanced perceptual quality. Despite higher sampling rates and channels, the quality of audio in training datasets remains inconsistent, often suffering from noise, dullness, and a lack of rhythm or structure. These problems, often reflected by the Mean Opinion Score (MOS), are rarely addressed. In terms of text-audio consistency, Make-an-audio 2 (Huang et al., 2023b) and WavCaps (Mei et al., 2024) have employed ChatGPT-assisted data augmentation to improve temporal relationships and accuracy in audio effect generation. Although studies like Music-llama (Liu et al., 2024) and LP-musiccaps (Doh et al., 2023) have introduced captioning approaches for music, few have explored the augmentation and utilization of synthetic data in large-scale music generation tasks.
@@ 4.1 Quality Information Injection
This process provides finer granularity of control within the following model and facilitates the ability of interpolative quality control during inference, enabling precise adjustments in $\mathbb{R}^{d}$.
@@ 5.2 Models and Hyperparameters
We utilize the LP-MusicCaps (Doh et al., 2023) caption model for ASM, FMA, and subsets of MTT and MSD that have weak or no captions. We use the official checkpoint from LAION-CLAP (Wu* et al., 2023)222music_speech_audioset_epoch_15_esc_89.98.pt for text-to-text and text-to-audio similarity calculations. Based on small scale subjective experientment, thresholds are set at $\rho_{1}=\rho_{2}=0.1$ to ensure any generated text or original caption not aligned well with the corresponding waveform is filtered out. Additionally, after filtering, generated text that fall below a threshold of $\rho_{3}=0.25$ are merged with original tags with the prompt: Merge this music caption “generated caption” with the ground truth tags “original tags”, and do not add any imaginary elements.. We use FLAN-T5-large (Peebles and Xie, 2023) as text encoder for all models.
@@ 5.3 Evaluation Metrics
For the human studies, we invite human raters to evaluate two aspects of the audio samples (i) overall quality (Ovl), and (ii) relevance to the text input (Rel). For the overall quality test, raters were asked to rate the perceptual quality of the provided samples with 1 to 5. For the text relevance test, raters were asked to rate the match between audio and text on a scale from 1 to 5. Our evaluators consist of people from various backgrounds, including professional music producers, video editors, staff responsible for the operations of music products and complete beginners who are unfamiliar with related knowledge. We evaluate randomly sampled file and each sample was evaluated by at least 5 raters.
@@ 6.2 Impact of Music Caption Refinement
Table 2: Objective evaluation results for music generation with diffusion based models and language model based models. We re-infer AudioLDM 2 and MusicGen with label †.
@@
MusicLM 1290M 280k 4.00 - - - - - - -
MusicGen † 1.5B 20k 3.80 1.22 - 0.31 5.38 1.01 1.92 0.18
Mousai 1042M 2.5k 7.50 1.59 - 0.23 - - - -
Jen-1 746M 5.0k 2.0 1.29 - 0.33 - - - -
AudioLDM 2 - Full 712M 17.9k 3.13 1.20 - - - - - -
AudioLDM 2 - Music † 712M 10.8k 4.04 1.46 2.67 0.34 2.77 0.84 1.91 0.28
@@
We conducted our ablation study on a subset of our training set, which includes ASM and FMA, totaling approximately 3,700 hours and 1.1 million clips. For evaluation, we utilized an out-of-domain set with 1,000 samples randomly selected from MTT (Law et al., 2009). Table  1 compares the model’s performance using different textual representations: sentences formed by merging original tags with commas ($\mathcal{D}_{o}$), generated captions ($\mathcal{D}_{g}$), and generated captions refined through filtering and fusion ($\mathcal{D}_{\text{merge}}$). During the filtering and fusion stage, 8.9% of the generated captions were filtered out, and 15.1% were fused with original tags using ChatGPT. Each model underwent training for 60,000 steps with a batch size of 64.
@@ 6.3 Compared with Previous Methods
We compared our proposed method with the following representative previous methods: AudioLDM 2 (Liu et al., 2023a), Mousai (Schneider et al., 2023) and Jen-1 (Li et al., 2023) which model music using spectral latent spaces, MusicLM (Agostinelli et al., 2023), and MusicGen Copet et al. (2024), which focus on modeling discrete representations.
@@ 7 Conclusion and discussion
In this study, we address the key challenges in the music generation domain, including model architecture design, large-scale uneven audio quality, and unaligned textual annotations, all of which impede the progress of TTM. By introducing a novel quality-aware training approach, along with MDT backbone and a three-stage text alignment scheme on large-scale open-source music dataset, we improve generation quality, musicality, and text alignment. In the future, we aim to further enhance and expand our model to achieve long-duration, high-sampling-rate, controllable, and highly interactive music generation. Additionally, we are committed to open-sourcing our model and checkpoints.
@@ A.3 Comparing with Using “Negative Prompt”
Based on the TableA.3, we found that any form of quality guidance improves the model’s generative performance. However, previous attempts to improve quality relied on the rare instances of ”low quality” in the dataset. This necessitated the careful design of numerous negative prompts to avoid generating low-quality results. Furthermore, the text embedding of ”low quality” might not be well disentangled during training, leading to an suboptimal results. Furthermore, while alternative guidance strategies (e.g. Autoguidance [Karras et al., 2024]) are capable of improving generation quality, they often come at the cost of significantly increased model parameters and sampling steps. In contrast, by incorporating trainable quality awareness, we are able to enhance overall generation quality without sacrificing sampling steps and inference memory consumption.

### v4

@@ Abstract
Text-to-music (TTM) generation, which converts textual descriptions into audio, opens up innovative avenues for multimedia creation. Achieving high quality and diversity in this process demands extensive, high-quality data, which are often scarce in available datasets. Most open-source datasets frequently suffer from issues like low-quality waveforms and low text-audio consistency, hindering the advancement of music generation models. To address these challenges, we propose a novel quality-aware training paradigm for generating high-quality, high-musicality music from large-scale, quality-imbalanced datasets. Additionally, by leveraging unique properties in the latent space of musical signals, we adapt and implement a masked diffusion transformer (MDT) model for the TTM task, showcasing its capacity for quality control and enhanced musicality. Furthermore, we introduce a three-stage caption refinement approach to address low-quality captions’ issue. Experiments show state-of-the-art (SOTA) performance on benchmark datasets including MusicCaps and the Song-Describer Dataset with both objective and subjective metrics. Demo audio samples are available at https://qa-mdt.github.io/, code and pretrained checkpoints are open-sourced at https://github.com/ivcylc/OpenMusic/.
@@ 1 Introduction
Figure 1: (a) The distribution curves of CLAP similarity and pseudo-MOS for large-scale open-source music databases AudioSet (Defferrard et al., 2016) and FMA (Defferrard et al., 2016), where darker areas represent higher text-audio consistency or audio quality. (b) Frechet Audio Distance (FAD) (Kilgour et al., 2018) scores on the MusicCaps test set obtained from models trained for 50K steps on AudioSet and FMA, using filter ratios of 0%, 33%, 66%, and 100% of low-quality data. Here, low-quality data is determined by a Pseudo-MOS score (Ragano et al., 2023) of less than 4.0. It can be inferred that performance consistently worsens with larger filter ratios.
@@
In the TTM field, high-quality music signals is scarce. This prevalent issue of low-quality data, highlighted in Figure 1 (a), manifests in two primary challenges. Firstly, most available music signals often suffer from distortion due to noise, low recording quality, or outdated recordings, resulting in diminished generated quality, as measured by pseudo-MOS scores from quality assessment models (Ragano et al., 2023). Secondly, there is a weak correlation between music signals and captions, characterized by missing, weak, or incorrect captions, leading to low text-audio similarity, which can be indicated by CLAP scores (Wu et al., 2023). These challenges, especially the inherent quality in the music signal itself, significantly hinder the training of high-performance music generation models, resulting in poor rhythm, noise, and inconsistencies with textual control conditions in the generated audio. Additionally, as shown in Figure 1 (b), directly filtering low-quality music, which robustly reduces the dataset size, leads to a consistent decline in model performance. Therefore, finding an effective training strategy for large-scale datasets with low-quality waveforms, mismatches, and missing labels has become an urgent challenge.
@@ 2 Related Work
Previous research has made efforts to improve the quality of generated audio, particularly in two key areas: waveform fidelity and the consistency between input text and generated content. Waveform quality can be compromised by issues like aliasing from low sampling rates and limited expressiveness due to monophonic representations, while models like MusicGen (Copet et al., 2024) and Stable Audio (Evans et al., 2024a, b), which directly model 32k and 44.1k stereo audio, have significantly enhanced perceptual quality. Despite higher sampling rates and channels, the quality of audio in training datasets remains inconsistent, often suffering from noise, dullness, and a lack of rhythm or structure. These problems, often reflected by the Mean Opinion Score (MOS), are rarely addressed. In terms of text-audio consistency, Make-an-audio 2 (Huang et al., 2023b) and WavCaps (Mei et al., 2024) have employed ChatGPT-assisted data augmentation to improve temporal relationships and accuracy in audio effect generation. Although studies like Music-llama (Liu et al., 2024) and LP-musiccaps (Doh et al., 2023) have introduced captioning approaches for music, few have explored the augmentation and utilization of synthetic data in large-scale music generation tasks. Additionally, Gui et al. (2024) overhaul FAD by introducing a sample-size–extrapolated score, curated high-quality reference sets, and music-aware embeddings, producing scores that align much more closely with human judgments of music quality.
@@ 4.1 Quality Information Injection
This process provides finer granularity of control within the following model and facilitates the ability of interpolative quality control during inference, enabling precise adjustments in $\mathbb{R}^{d}$. In later stages, the quality embedding is treated as a token on par with every latent audio patch, participating in the attention computations to enable interaction.
@@ 5.2 Models and Hyperparameters
We utilize the LP-MusicCaps (Doh et al., 2023) caption model for ASM, FMA, and subsets of MTT and MSD that have weak or no captions. We use the official checkpoint from LAION-CLAP (Wu et al., 2023)222music_speech_audioset_epoch_15_esc_89.98.pt for text-to-text and text-to-audio similarity calculations. Based on small scale subjective experienment, thresholds are set at $\rho_{1}=\rho_{2}=0.1$ to ensure any generated text or original caption not aligned well with the corresponding waveform is filtered out. Additionally, after filtering, generated text that fall below a threshold of $\rho_{3}=0.25$ are merged with original tags with the prompt: Merge this music caption “generated caption” with the ground truth tags “original tags”, and do not add any imaginary elements.. We use FLAN-T5-large (Peebles and Xie, 2023) as text encoder for all models.
@@ 6.2 Impact of Music Caption Refinement
MusicLM 1290M 280k 4.00
MusicGen 1.5B 20k 3.80 1.22 0.31 5.38 1.01 1.92 0.18
Mousai 1042M 2.5k 7.50 1.59 0.23
Jen-1 746M 5.0k 2.00 1.29 0.33
AudioLDM 2 – Full 712M 17.9k 3.13 1.20
AudioLDM 2 – Music 712M 10.8k 4.04 1.46 2.67 0.34 2.77 0.84 1.91 0.28
@@
Table 2: Objective evaluation results for music generation with diffusion-based and language-model-based approaches. Methods we re-inferred are marked with †.
@@
We conducted our ablation study on a subset of our training set, which includes ASM and FMA, totaling approximately 3,700 hours and 1.1 million clips. For evaluation, we utilized an out-of-domain set with 1,000 samples randomly selected from MTT (Law et al., 2009). Table  2 compares the model’s performance using different textual representations: sentences formed by merging original tags with commas ($\mathcal{D}_{o}$), generated captions ($\mathcal{D}_{g}$), and generated captions refined through filtering and fusion ($\mathcal{D}_{\text{merge}}$). During the filtering and fusion stage, 8.9% of the generated captions were filtered out, and 15.1% were fused with original tags using ChatGPT. Each model underwent training for 60,000 steps with a batch size of 64.
@@ 6.3 Compared with Previous Methods
We compared our proposed method with the following representative previous methods: AudioLDM 2 (Liu et al., 2023a), Mousai (Schneider et al., 2023) and Jen-1 (Li et al., 2023) which model music using spectral latent spaces, MusicLM (Agostinelli et al., 2023), and MusicGen (Copet et al., 2024), which focus on modeling discrete representations.
@@ 7 Conclusion and Discussion
In this study, we address the key challenges in the music generation domain, including model architecture design, large-scale uneven audio quality, and unaligned textual annotations, all of which impede the progress of TTM with quality, musicality, and text alignment. In the future, we aim to further enhance and expand our model to achieve long-duration, high-sampling-rate, controllable, and highly interactive music generation.
@@ Acknowledgements
Statement
Chang Li and Ruoyu Wang contributed equally. Jun Du supervised the entire project and serves as the corresponding author. Further discussion and statement can be found in the Appendix Supplementary Statement.
@@ A.2 Patchify Strategy
Supplementary Note. We employ patches with a highly unbalanced aspect ratio because, in audio spectrograms, the time axis is typically much longer than the frequency axis and Table A.2 shows no significant performance gain when the temporal granularity is made finer. Furthermore, the perceived decline in melody quality hinges strongly on localized subjective judgments, and it remains unclear whether this phenomenon would differ with larger-scale models and datasets.
@@ A.3 Comparing with Using “Negative Prompt”
Based on the TableA.3, we found that any form of quality guidance improves the model’s generative performance. However, previous attempts to improve quality relied on the rare instances of ”low quality” in the dataset. This necessitated the careful design of numerous negative prompts to avoid generating low-quality results. Furthermore, the text embedding of ”low quality” might not be well disentangled during training, leading to an suboptimal results. Furthermore, while alternative guidance strategies (e.g. Autoguidance [Karras et al., 2024]) are capable of improving generation quality, they often come at the cost of significantly increased model parameters and sampling steps. In contrast, by incorporating trainable quality awareness, we are able to enhance overall generation quality without sacrificing sampling steps and inference memory consumption. We also leave combining training-based quality-aware strategy and negative-guidance strategy as future work.
Compared with negative-prompt guidance, we believe our approach offers the following key advantages:
  • As shown in our main experiments, injecting quality information directly at the token-embedding level yields much stronger and more precise control than injecting it through the text encoder’s prompt.
  • Negative-prompt guidance requires crafting accurate negative descriptions, which increases the burden of prompt engineering and the effort needed to encode specific quality cues, thereby limiting its scalability to other domains and tasks.
  • When quality attributes are expressed in natural language, they inevitably interact inside the text encoder with other audio-description tokens; this cross-talk between otherwise independent features can lead to sub-optimal outputs.
Supplementary Statement
Scope of the p-MOS Predictor.
The perceptual MOS (p-MOS) scores reported in this study are derived from a music-quality prediction model whose training distribution—owing to the experiment’s temporal constraints—is most closely aligned with vinyl-record characteristics and restricted at 16 kHz. Future work should incorporate a broader, more finely differentiated set of evaluative dimensions that span overall musical quality, musicality, signal fidelity, intrinsic timbre, and general perceptual preference. Expanding both the taxonomic granularity and the pool of human-annotated samples will enable a more accurate, widely applicable quality-assessment framework [Tjandra et al., 2025].
Sample Length, Commercial Baselines, and Generality of Our Methods.
Because open-source academic music datasets are typically short, we trained and evaluated our system on 10-second excerpts and did not benchmark against cutting-edge commercial generators such as Suno [Suno AI, 2024] or Udio [Udio, 2024], which can synthesize full-length tracks. Although our U-Net variant naturally scales to arbitrarily long outputs—and diffusion-trajectory fusion can extend generation without retraining [Bar-Tal et al., 2023; Dai et al., 2025]—we emphasize that our proposed quality-oriented strategies and architectural insights are generation-paradigm agnostic. They can benefit future systems regardless of training regimen, sequence length, channel configuration, sampling rate, or compression constraints. While training on 10-second excerpts provides a practical proxy for validating algorithmic performance, the music modality more readily captures subjective qualities such as melodic appeal and perceived listening quality.
Data Ethics and Annotation Pipeline.
All training audio was curated to exclude copyright-restricted material, and large-language-model APIs were employed solely for caption refinement under fair-use terms. Nevertheless, more recent multimodal annotators—e.g., Gemini [Team et al., 2023], Kimi-Audio [Ding et al., 2025], or Audio Flamingo [Kong et al., 2024; Ghosh et al., 2025]—offer richer contextual labeling; integrating them to enhance caption fidelity and diversity is left for future work.